### v1

# Direct Preference Knowledge Distillation for Large Language Models

Abstract
In the field of large language models (LLMs), Knowledge Distillation (KD) is a critical technique for transferring capabilities from teacher models to student models. However, existing KD methods face limitations and challenges in distillation of LLMs, including efficiency and insufficient measurement capabilities of traditional KL divergence. It is shown that LLMs can serve as an implicit reward function, which we define as a supplement to KL divergence. In this work, we propose Direct Preference Knowledge Distillation (DPKD) for LLMs. DPKD utilizes distribution divergence to represent the preference loss and implicit reward function. We re-formulate KD of LLMs into two stages: first optimizing and objective consisting of implicit reward and reverse KL divergence and then improving the preference probability of teacher outputs over student outputs. We conducted experiments and analysis on various datasets with LLM parameters ranging from 120M to 13B and demonstrate the broad applicability and effectiveness of our DPKD approach. Meanwhile, we prove the value and effectiveness of the introduced implicit reward and output preference in KD through experiments and theoretical analysis. The DPKD method outperforms the baseline method in both output response precision and exact match percentage. Code and data are available at https://aka.ms/dpkd.
Direct Preference Knowledge Distillation for Large Language Models
Yixing Li1thanks:  Contribution during internship at Microsoft Research., Yuxian Gu211footnotemark: 1, Li Dong3, Dequan Wang1, Yu Cheng4, Furu Wei 3 1Shanghai Jiao Tong University   2Tsinghua University 3Microsoft Research   4The Chinese University of Hong Kong {lyxing0, dequanwang}@sjtu.edu.cn   guyx21@mails.tsinghua.edu.cn {lidong1,fuwei}@microsoft.com   chengyu@cse.cuhk.edu.hk
1 Introduction
In the era of Large language models (LLMs), a series of models and techniques have demonstrated great capabilities Zhang et al. (2022a); Chowdhery et al. (2023) , and their powerful capabilities are usually ascribed to the increase in the size of training data and model scale Kaplan et al. (2020); Anil et al. (2023). However, the increase in the size of large models also brings expensive computing costs and difficulties Hoffmann et al. (2022). How to maintain performance and conduct efficient training while increasing the scale of language models has become an important issue. Knowledge distillation (KD; Hinton et al., 2015) trains a relatively compact student model by simulating the output distribution and behavior of the teacher model, which is an effective method under limited computing resources. For white-box models, we can have access to richer information such as the distribution of logits of the model Gou et al. (2021). This provides us with more value because the distribution of the hidden layer represent the underlying knowledge of LLMs, leading to better effect and output in knowledge transfer Wang et al. (2020).
Recent work has recognized the shortcomings of KL divergence in traditional KD in which of LLMs Gu et al. (2023), and explored other forms of KL divergence like reverse KL divergence. Work on KL divergence Wu et al. (2024) proposed different KLD metrics for different stages of their respective shortcomings, and proposed improvement methods to expand KL divergence into flexible distance form. In fact, KL divergence is insufficient under the condition of a stronger teacher model Huang et al. (2022); Cho and Hariharan (2019); Son et al. (2021), requiring the introduction of additional objectives and innovative knowledge distillation procedures. We start from another perspective and consider a novel optimization objective to compensate for the above problems while maintaining high efficiency.
In this work, we propose Direct Preference Knowledge Distillation (DPKD) for knowledge distillation of LLMs. It is shown that LLMs can serve as an implicit reward function Yuan et al. (2024); Rafailov et al. (2024b). Due to the deficiency of KL divergence Gu et al. (2023); Huang et al. (2022), we define implicit reward function as a supplement to the KL distance. We re-formulate the KD of white-box LLMs as follows, first maximize the optimization function consisting of implicit reward and reverse KL divergence, and then improve the preference probability of teacher outputs over student outputs. These settings compensate for the shortcomings of KL divergence while preserving high training efficiency. We derive the training objective to obtain a concise final form, and prove the role of the reward and preference form we introduced in KD through theoretical derivation.
Finally, we conducted experiments on our DPKD method in the instruction tuning topic. We use different families of LLMs including GPT-2 Radford et al. (2019) and OPT Zhang et al. (2022b) with parameter sizes ranging from 120M to 13B and multiple datasets to validate our method. We evaluate the results with RougeL Lin (2004) metric , which is commonly used in instruction fine-tuning tasks. Our results show that the DPKD method outperforms the baseline method and has advantages over a wide range of generation lengths. We also conducted experiments on the implicit reward function in the reformulation of distillation to verify effectiveness, and experiments on the forms of preference to show the potential of preference modeling in KD and provide ideas for subsequent research. Our contributions are thus three-fold:
  • We propose a novel framework called Direct Preference Knowledge Distillation (DPKD), which provides a new perspective besides designing different KL divergences for knowledge distillation of large models.
  • We compare DPKD to five baseline methods across two commonly used LLMs (GPT-2 and OPT) and three datasets. We also provide a detailed derivation and theoretical analysis to demonstrate the effectiveness of the reward and preference model introduced in this paper.
  • We provide additional experiments on other preference objectives and observations on implicit reward. We illustrate the effectiveness and potential of our reformulation of the KD process, providing inspiration for subsequent work.
2 Methods
2.1 Preliminaries
2.1.1 Sequence-Level Knowledge Distillation
[…]
2.1.2 Preference Model
Given two comparable objects or events, a common model for comparing the probabilities of their selection is the Bradley-Terry (BT; Bradley and Terry, 1952) model. We consider output $y_{1}$ and $y_{2}$, and the reward function that measures the gain of choosing an output is denoted $r(y)$. BT model can be used to measure the probability that we choose $y_{1}$ instead of $y_{2}$:
$Pr\left(y_{1}\succ y_{2}\right)$ $=\frac{\exp\left(r\left(y_{1}\right)\right)}{\exp\left(r\left(y_{1}\right)\right)+\exp\left(r\left(y_{2}\right)\right)}$ (1)
$=\sigma\left[r\left(y_{1}\right)-r\left(y_{2}\right)\right],$
where $\sigma$ is the function $\sigma(x)\mathrel{\coloneqq}\frac{1}{1+\exp(-x)}$.
2.2 DPKD: Direct Preference Knowledge Distillation
In the context of knowledge distillation of LLMs, the distribution difference measured by KL divergence is usually considered to be the only criterion for the distillation target. But KL divergence is not enough. From the perspective of measuring distributions, KL divergence is 0 if and only if $q_{\theta}$ and $p$ are the same distribution in the case of discrete variables, or almost identical in the case of continuous variables. KL divergence measures the difference between two distributions, but it is not really a distance because it is asymmetric Manning and Schutze (1999) . Therefore, some researchers have improved the KL divergence by combining two symmetrical items (Jensen-Shannon divergence Nielsen (2021)). In the field of KD, researchers have explored the inadequacy of KL divergence in KD Cho and Hariharan (2019); Mirzadeh et al. (2020). Some works Gu et al. (2023); Wu et al. (2024) hope to redesign KL divergence, and some worksHuang et al. (2022) add new objectives besides KL divergence.
Based on the above considerations, we denote the implicit reward $r_{p}\left(\mathbf{y}|\mathbf{x}\right)$ as a supplement to the KL divergence in the optimization objective. We formulate the optimization goal as:
$\max_{\theta}\mathbb{E}\,[\,r_{p}(y|x)-\beta\,\text{KLD}\,\,(q_{\theta}(y|x))\|p(y|x))\,].$ (2)
We can get the optimal solution of Equation 2 through the following steps. Firstly transform Equation 2 as follows:
$\min_{\theta}\mathbb{E}\left[\log\frac{q_{\theta}(y\mid x)}{p(y\mid x)}-\frac{1}{\beta}r(x,y)\right]$ (3)
$=$ $\min_{\theta}\mathbb{E}\left[\log\frac{q_{\theta}(y\mid x)}{q^{*}(y\mid x)}-\log Z(x)\right],$
where $q^{*}(y|x)\mathrel{\coloneqq}\frac{1}{Z(x)}p(y|x)\exp\left(\frac{1}{\beta}r(x,y)\right)$, and $Z\left(x\right)$ is the normalized function of the distribution, which is required to be independent of $\theta$ and $y$. It is defined as $Z(x)\mathrel{\coloneqq}\sum_{y}p(y|x)\exp\left(\frac{1}{\beta}r(x,y)\right)$. Following the work, we can derive the optimal solution of Equation 3:
$q_{\theta}(y\mid x)=q^{*}(y\mid x).$ (4)
From Equation 4 we can obtain that:
$r^{*}(x,y)=\beta\log\frac{q^{*}(y\mid x)}{p(y\mid x)}+\beta\log Z(x).$ (5)
Given the same prompt x, the outputs of the student model and the teacher’s model are respectively denoted as $y_{s}$ and $y_{t}$. The purpose of KD is to fit the distribution of the student model to teacher model, which can also be understood as we expect the student model to have a greater probability of outputting results similar to the teacher model. From the BT model, we can obtain the following probability:
$p$ ${}^{*}\left(y_{t}\succ y_{s}\mid x\right)$ (6)
$=\sigma\left(\beta\log\frac{q_{\theta}\left(y_{t}\mid x\right)}{p\left(y_{t}\mid x\right)}-\beta\log\frac{q_{\theta}\left(y_{s}\mid x\right)}{p\left(y_{s}\mid x\right)}\right).$
The complete derivation of the above procedure is in the Appendix A. Since our goal is to maximize the probability of model outputing $y_{t}$ rather than $y_{s}$, which is $\max_{\theta}\,\,\mathbb{E}\log\,p^{*}\left(y_{t}\succ y_{s}\mid x\right)$. It is formulated as:
$\min_{\theta}-\mathbb{E}\left[\log\sigma\left(\beta\log\frac{q_{\theta}\left(y_{t}\right)}{p\left(y_{t}\right)}-\beta\log\frac{q_{\theta}\left(y_{s}\right)}{p\left(y_{s}\right)}\right)\right].$ (7)
where $q_{\theta}\left(y_{t}\right)$ denotes $q_{\theta}\left(y_{t}\mid x\right)$ for short, and same for $p$ and $y_{s}$.
2.2.1 Learning Objective
Following the framework derivation of DPKD, the loss is defined as:
$\mathcal{L}=-\mathbb{E}\left[\log\sigma\left(\beta\log\frac{q_{\theta}\left(y_{t}\right)}{p\left(y_{t}\right)}-\beta\log\frac{q_{\theta}\left(y_{s}\right)}{p\left(y_{s}\right)}\right)\right].$ (8)
It is shown by works Meng et al. (2024); Gu et al. (2023) that models tend to introduce bias and produce short responses without length normalization. We add the length normalization factor to the distillation loss as follows:
$\log\sigma\left(\frac{\beta}{\lvert y_{t}\rvert}\log\frac{q_{\theta}\left(y_{t}\right)}{p\left(y_{t}\right)}-\frac{\beta}{\lvert y_{s}\rvert}\log\frac{q_{\theta}\left(y_{s}\right)}{p\left(y_{s}\right)}\right).$ (9)
Following work Ouyang et al. (2022); Gu et al. (2023), we add the language modeling loss in order to preserve performance on canonical NLP benchmarks. The complete algorithm process is shown in Algorithm 1.
Algorithm 1 Direct Preference Knowledge Distillation (DPKD)
0:    Instruction tuning datasets $D$; Pre-training corpus $D_{p}$; Fine-tuned teacher model with distribution $p$;Initialized student model conducted SFT on $D_{p}$ with outpu distribution $q_{\theta}$;
0:    Student model parameter conducted distillation training;
1:  for epoch in epochs do
2:     for batch in datasets $D$ and $D_{p}$ do
3:        Compute responses from teacher and student model and obtain $y_{t}$ and $y_{s}$;
4:        Computes four log items $\log p(y_{t})$, $\log p(y_{s})$, $\log q_{\theta}(y_{t})$ and $\log q_{\theta}(y_{s})$
5:        Compute distill loss $\mathcal{L}_{kd}=-\sum_{\left(x\sim\mathcal{D}\right)}$ $\log\sigma\left(\frac{\beta}{\lvert y_{t}\rvert}\log\frac{q_{\theta}\left(y_{t}\right)}{p\left(y_{t}\right)}-\frac{\beta}{\lvert y_{s}\rvert}\log\frac{q_{\theta}\left(y_{s}\right)}{p\left(y_{s}\right)}\right)$
6:        Compute language modeling loss $\mathcal{L}_{pt}=-\sum_{d\sim\mathcal{D}_{p}}\log q_{\theta}(d)$
7:        Compute loss $\mathcal{L}=\mathcal{L}_{kd}+\lambda\cdot\mathcal{L}_{pt}$ and gradient $\nabla_{\theta}\mathcal{L}$
8:        Gradient Update $\theta^{(t+1)}=\theta^{(t)}-\alpha\nabla_{\theta}\mathcal{L}$
9:     end for
10:  end for
11:  return  Student model parameter $\theta$;
3 Analysis
3.1 Gradient Derivation
Based on our target definition, we can derive the gradient of DPKD through the basic chain rule and perform analysis similar to work Rafailov et al. (2024b). From Equation 8, we can perform variable substitution and define $u\mathrel{\coloneqq}\beta\log\frac{q_{\theta}\left(y_{s}\mid x\right)}{p\left(y_{s}\mid x\right)}-\beta\log\frac{q_{\theta}\left(y_{t}\mid x\right)}{p\left(y_{t}\mid x\right)}$, and the gradient of the loss can be expressed as
$\nabla_{\theta}\mathcal{L}$ $=-\nabla_{\theta}\mathbb{E}_{\left(x,y_{t},y_{s}\right)\sim\mathcal{D}}\left[\log\sigma\left(u\right)\right]$ (10)
$=-\mathbb{E}_{\left(x,y_{t},y_{s}\right)\sim\mathcal{D}}\left[\frac{\sigma^{\prime}(u)}{\sigma(u)}\nabla_{\theta}(u)\right].$
We can derive the complete form of gradient:
$\nabla_{\theta}\mathcal{L}=$ $-\mathbb{E}\left[\beta\sigma\left(\beta\log\frac{q_{\theta}\left(y_{t}\right)}{p\left(y_{t}\right)}-\beta\log\frac{q_{\theta}\left(y_{s}\right)}{p\left(y_{s}\right)}\right)\right.$ (11)
$\left.\left[\nabla_{\theta}\log q_{\theta}\left(y_{t}\right)-\nabla_{\theta}\log q_{\theta}\left(y_{s}\right)\right]\right].$
The full derivation of the gradient is in the Appendix B.
3.2 Optimizing DPKD Is Optimizing The Q Function
Following work Rafailov et al. (2024a), we conduct a theoretical analysis of DPKD and relate the reward function in KD we introduced to the Q function..We consider the sequence generation task, and the data set $\mathcal{D}=\{(\mathbf{x},\mathbf{y})\}^{N}$, where the prompt is denoted as $\mathbf{x}=\{x_{0},\cdots,x_{l-1}\}$. We denote the vocabulary as $\mathbf{V}$, each word $x$ is included in the vocabulary $\mathbf{V}$. Given input prompt $\mathbf{x}$, the model output is $\mathbf{y}=\{y_{0},\cdots,y_{m-1}\}$, where $m$ is the maximum generated length. At each time step, the generated $y_{t}$ is conditionally generated based on the generated sequence $\{\mathbf{x},\mathbf{y_{t-1}}\}$.
From another perspective of Markov Decision Process (MDP), the process of sequence generation by the model is regarded as the generation process of a Markov decision chain. The state is denoted as $s_{t}=\{x_{0},\ldots,x_{l-1},y_{0},\ldots,y_{t-1}\}$, and the action is $y_{t}\in\mathbf{V}$, which is chosen based on generated sequence $s_{t}$. The transition function from the current state to the next state is the LLMs we selected.
We analyze the implicit reward function of DPKD from $Q$ function, which is a perspective in reinforcement learning. $Q$ function is specifically explained as follows: $Q\left(s,a\right)$ refers to the expectation of the benefit obtained by taking action $a$ (a word selected from the vocabulary, obtained by taking the maximum from the output distribution or sampling by top-k) in state $s$ (generated sequence $\{x,y\}$). The general represent of $Q$ is:
$Q^{\theta}(s_{t},a_{t})=\mathbb{E}[R_{t+1}+\gamma R_{t+2}+...\mid s_{t},a_{t}].$ (12)
Following the MDP perspective of sequence generation and according to the work , the fixed point solution of Equation 12 is:
$q_{\theta}^{*}(\mathbf{a}_{t}|\mathbf{s}_{t})=e^{(Q^{*}(\mathbf{s}_{t},\mathbf{a}_{t})-V^{*}(\mathbf{s}_{t}))/\beta}.$ (13)
Any valid $Q$ function needs to satisfy the Belmman equation, from which we can write the current step $Q^{*}$ function for the optimal strategy (that is, the most optimal student model parameters) and the reward function $r$ as:
$Q^{*}(\mathbf{s}_{t},\mathbf{a}_{t})=r(\mathbf{s}_{t},\mathbf{a}_{t})+\beta\log p(\mathbf{a}_{t}|\mathbf{s}_{t})+V^{*}(\mathbf{s}_{t+1}),$ (14)
where $V^{*}(\mathbf{s})$ is defined to be zero if $s$ is the end of sequence (EOS). Following the work , we can use the $Q$ function instead of $r$ and substitute it into the Bellman equation.
Now we can use the $Q$ function to re-derive the DPKD method. We use the $Q$ function and the $V$ function to represent the sum of the reward function written at the current time step t. By transforming the Equation 14 and summing over time t, and substitute Equation 13 into the resulting sum, we could obtain:
[…]
According to the multi-step generalization of the Bradley-Terry preference model, namely the Plackett-Luce model, and substituting the Equation 15 into it, we can obtain:
$p(\tau^{t}\succeq\tau^{s})=\sigma\left(\beta\log\frac{q_{\theta}\left(y_{t}\right)}{p\left(y_{t}\right)}-\beta\log\frac{q_{\theta}\left(y_{s}\right)}{p\left(y_{s}\right)}\right),$ (16)
where $\tau$ refers to the sequence trajectory generated by the model, and in the sequence generation task of the large model, it refers to the generated text. Replace $\tau^{t}$ and $\tau^{s}$ in the formula with $y_{t}$ and $y_{s}$ respectively, and the above formula is exactly the same as Equation 6. Thus we get the complete process of deriving the DPKD of this work from the $Q$ function. Therefore, we can link the DPKD method of this work with the concept of $Q$ function in reinforcement learning and Markov decision chain.
3.3 Your Language Model Is Secretly Implicit Reward Function During KD
From the experimental results, we implicitly optimized a reward function in the KD process by redefining the formula and objectives of KD, and our experimental results show the role of the reward function in the KD training process. This shows that the KD in our experiment is actually optimizing a reward function, which is an unobserved perspective.
From the perspective of theoretical analysis, the reward function not only serves as a weight term in the gradient of our DPKD, but also as a hub connecting our DPKD method and the Q function of reinforcement learning.
Furthermore, from the results in Section 4.4, we can see that experiments with different preference expressions related to reward functions show quite different results, and some preference expressions even show promising results, even though we did not do much parameter fine-tuning. This shows that using reward functions to express preferences in the knowledge distillation process is effective, and there is still a lot of value to explore, which also provides ideas for subsequent work.
4 Experiments
4.1 Experiments Setup
Tasks and Metrics.
We conduct experiments and analysis on the task of instruction tuning. The topics of instruction tuning range from summarizing to completing the request, and the model is required to complete the response according to instruction, prompt, and input. We specifically select data with instructions and outputs longer than 10 words in the dataset to ensure that the model aims to generate high-quality results that are long enough. The RougeL Lin (2004) score is used to measure the quality of model generation, as Wang et al. (2022) has shown that RougeL is suitable for large-scale instruction tuning evaluation.
Datasets.
For the datasets, we use the following datasets: (1) Dolly Conover et al. (2023) including $\sim$15k instruction/response generated in capability domains from the InstructGPT, (2) Self-Inst Wang et al. (2023) consisting of 252 expert-written tasks and their instructions motivated by user-oriented applications, (3) SuperNatural-Instructions (S-NI; Wang et al., 2022) containing over 1.5k tasks and covering 76 distinct task types. We filter the data whose length exceeds the model processing length and select the part with response length $\left[11,\infty\right]$ as training data to keep the same settings as the validation set and test machine.
Models.
For teacher models, we select GPT-2 Radford et al. (2019) (1.5B) and OPT Zhang et al. (2022b) (13B). The student models are GPT-2 with 120M and 340M parameters and OPT with 1.3B parameters respectively. These models have been pre-trained on the $D_{p}$ datasets.
Baselines.
The baseline methods we selected include: (1) Supervised Fine-Tuning (SFT) directly fine-tune on labels of dataset; (2) KD Hinton et al. (2015) is also called word-level KD, which uses the teacher’s output distributions as supervision; (3) Sequence-level KD (SeqKDKim and Rush, 2016; Zhou et al., 2024) directly distills the student model on the data generated by the teacher model; (4) MiniLLM Gu et al. (2023) leverages reverse KL divergence to perform distillation on LLMs ;(5) Adaptive Kullback-Leiber (AKLWu et al., 2024) uses a combination of forward and reverse KL with dynamic weights to perform distillation.
4.2 Results
Model #Param Method DollyEval SelfInst S-NI
GPT-2 1.5B Teacher 27.6 14.3 27.6
120M SFT 23.3 10.0 16.3
KD 22.8 10.8 13.4
SeqKD 22.7 10.1 16.4
MiniLLM 24.6 13.2 25.3
AKL 23.9 - 19.2
DPKD (ours) 24.6 13.8 25.4
GPT-2 1.5B Teacher 27.6 14.3 27.6
340M SFT 25.5 13.0 25.1
KD 25.0 12.0 23.7
SeqKD 25.3 12.6 22.9
MiniLLM 25.4 15.6 27.4
DPKD (ours) 26.0 14.3 27.6
OPT 13B Teacher 29.2 18.4 30.4
1.3B SFT 26.0 11.4 23.1
KD 25.4 12.2 21.9
SeqKD 26.1 12.7 21.4
MiniLLM 26.7 14.8 28.6
DPKD (ours) 27.2 14.9 28.7
Table 1: Experimental results with GPT-2 and GPT model families. The teacher models of the GPT-2 and OPT are 1.5B and 13B respectively. The student model ranges from 120M to 1.3B. The scores are RougeL scores, and the experiments cover the DollyEval, SelfInst, and S-NI datasets. All experiments are conducted with the same seed. AKL results are taken from Wu et al. (2024)
We conducted experiments based on the Algorithm 1 and the experimental settings in Section 4.1, and the experimental results are shown in Table 1. In this section, we conduct a preliminary analysis of the results including the RougeL score, the GPT-4 score, and an example of the implicit reward function introduced in this article.
From the experimental results we could see that our method shows great performance compared with the various baseline method. We conduct KD experiments on datasets with different distributions and LLMs of different sizes, showing that our method is applicable to a wide range of models and data and retains good performance. In some specific datasets, the student model trained by our knowledge distillation method is even close to the performance of the teacher model.
Figure 1: GPT-4 evaluation of different methods. Our DPKD method outperforms other baselines and is closest to the reference responses.
We collected the output texts of DPKD and the baseline method of GPT-2 Base, and scored them together with the labels of the dataset by GPT-4 to evaluate the quality of the instruction tracking generated text of each method compared with the label. Due to the input context text restrictions of GPT-4, we used the generation results and labels of the first 100 datasets of DollyEval. Although the output results of each method are still a certain distance away from the label, the generation results of the method proposed in this paper are close to the evaluation of the label, which is higher than baselines.
4.3 Reward Observation
Figure 2: We report the reverse KL divergence, and estimated implicit reward during training. The color of points represents the training epochs. The end of the training falls in the upper left corner, where the KL divergence is low and the reward is high.
We analyzes the implicit reward function involved in this paper, and analyzes the role of the reward function introduced in KD from both theoretical and experimental perspectives. According to the derivation of the gradient in Section 3.1, the reward function plays the role of a weight in the gradient update process. When the model parameter’s estimate of the reward function indicates that the current output is biased towards error, the value of the weight will be larger, and thus the gradient will be larger. When the current performance of the model is already better and the output distributions of the student model and the teacher model are close, the weight term we mentioned earlier will be low, so the model will tend to converge.
We hope to examine the trend of KL divergence and reward function of DPKD. We use an unbiased approximation to the reward function, which is $\hat{r}^{*}(x,y)=\beta\log\frac{q_{\theta}(y|x)}{p(y|x)}$ Based on this and the definition of reward estimation, we plot the two variables of reward function and KL divergence for the same model (here we use the GPT-2 Base model) at different training periods. We can get the following results.
Figure 3: Reward, KLD, and reverse KLD curves during the distillation process of GPT-2 Base. KLD and reverse KLD show similar trends.
We visualize the trend of reward during training and the reverse KL divergence distance between the student and teacher models over training epochs in Figure 2. From the visualization results, we can see that although some model checkpoints are very close to the KL divergence of the teacher model (all within the range of $0.04\pm 0.02$), the difference in the reward function value still causes the difference in RougeL scores. From the curve trend in the figure, it can be seen that the optimal point of model training is taken at a left-point arc vertex, that is, the student model has the lowest KL divergence from the teacher and the reward function is at a higher value. The size of the reward function may be related to the hyperparameter $\beta$ in training, which is used to adjust the relative proportion of the KL divergence and the reward function.
Figure 3 shows the changes in the implicit reward function, KLD, and rKLD during training. We can see that the estimated implicit reward gradually increases within a range. The trends of KLD and rKLD are roughly the same, but the specific values are slightly different, which is consistent with the conclusions of previous work Gu et al. (2023); Wu et al. (2024).
4.4 Variants of Preference Objective
In the subject of fine-tuning language models with human feedback, there are many optimization methods based on preference. These different optimization methods are all based on the original work DPO, and modify the preference form, normalization method, etc. on this basis. In the field of knowledge distillation of large models, the function form derived in this paper is a relatively basic form. So similarly, we also conducted preliminary experiments on different forms of preference calculation methods, and hope to bring inspiration to subsequent research. The experiments conducted in this chapter did not perform special parameter fine-tuning.
Preference Form RougeL
IPO Azar et al. (2024) 18.78
CPO Xu et al. (2024) 24.31
SimPO Meng et al. (2024) 24.46
DPKD (Ours) 24.62
Table 2: Variants of preference objective. The definition of each preference is as follows: (1) SimPO = $-\log\sigma\left(\frac{\beta}{\left|y_{t}\right|}\log q_{\theta}\left(y_{t}|x\right)-\frac{\beta}{\left|y_{s}\right|}\log q_{\theta}\left(y_{s}|x\right)-\gamma\right)$;
(2) CPO = $-\log\sigma\left(\beta\log\frac{q_{\theta}\left(y_{t}|x\right)}{q_{\theta}\left(y_{t}|x\right)}\right)-\log q_{\theta}\left(y_{t}|x\right)$;
(3) IPO = $\left(\log\frac{q_{\theta}\left(y_{t}|x\right)}{p\left(y_{t}|x\right)}-\log\frac{q_{\theta}\left(y_{s}|x\right)}{p\left(y_{s}|x\right)}-\frac{1}{2\tau}\right)^{2}$.
We experimented with four forms of preference: IPO, CPO, SimPO. Their definitions and experimental results are shown in Figure 2.
From the experimental results, we could see that although other forms of preference do not exceed our method, they also show certain effectiveness. Pal et al. (2024) showed that in the context of learning from preference data, the form of the preference also affects how the relative scores of $y_{t}$ and $y_{s}$ grow during the training process. Our results show that these problems and solutions are also applicable to the field of knowledge distillation, which also provides inspiration for future work.
4.5 Ablation Studies
Setting RougeL $\uparrow$ rKLD $\downarrow$ Reward $\uparrow$
DPKD 24.62 0.07 0.97
  w/o LM Loss 23.47 0.24 0.93
  w/o Length Norm. 22.79 0.23 0.90
Table 3: Ablation studies of DPKD. “rKLD” (reverse KL divergence) is the minimum value of the student distance from the teacher model during training. “Reward” is the maximum value. LM loss represents the language modeling loss introduced in Section 2.2.1
We conduct ablation experiments on the experimental setting and measure the scores on the GPT-2 Base model and DollyEval. We evaluate the results of removing length normalization and language modeling loss. In Table 3, KLD and rKLD represent the minimum values of KL divergence and reverse KL divergence from the teacher model during training, respectively. To ensure consistent distribution conditions, the distribution difference is calculated based on the first token output. The reward is the maximum value during training and is calculated in the same way as above. From the results, we can see that the components added in the experiment have a positive effect on the results, especially the improvement of implicit reward. The complete method performs well with the teacher model under both reverse KL and forward KL metrics, and has the highest implicit reward and score.
4.6 Results of Different Generation Lengths
Figure 4: RougeL score with different ranges of generation lengths. In the case of different ranges of reference label lengths, DPKD scores higher than the baseline. In particular, DPKD stands out when the golden response length is in the middle range. The raw RougeL scores of each method are provided in the Appendix C.
Method $\left(0,30\right)$ $\left(30,70\right)$ $\left(70,\infty\right)$
SeqKD 5.15 0.00 0.00
MiniLLM 5.30 0.93 0.63
DPKD 5.57 1.85 0.63
Table 4: Exact match with different ranges of generation lengths. Our method outperforms the baselines in different ranges of reference lengths, and the results of other baselines tend to be 0 when the length of the golden response is longer.
We experiment on subsets of the test set with different response lengths to test the effectiveness of our approach for scenarios with different output length requirements. We divide the DollyEval test set into three subsets according to the length of the golden response: $\left(0,30\right)$, $\left(30,70\right)$ and $\left(70,\infty\right)$, where 30 and 70 are the median and mean of the response length, respectively. We then evaluate the response results of different baselines and DPKD methods on these subsets.
The results of RougeL and exact match are shown in Figure 4 and Table 4. Our DPKD method performs well on all subsets of response length splits. In particular, DPKD exceeds the baseline by the most in the results with response lengths in the middle. In the exact match results, the scores of many baseline methods drop to 0 when the golden response length is longer, while DPKD still performs well and exceeds the baseline.
5 Related Work
Knowledge Distillation (KD)
is a widely used technique in model training and fine-tuning, and was first proposed in the work Bucila et al. (2006), Hinton et al. (2015) further enriched and conducted studies on KD. Researchers have explored the application of KD in text generation tasks Song et al. (2020); Zhang et al. (2023); Jiao et al. (2020); Sun et al. (2019). The standard form of KD in NLP is to minimize the KLD between student model and teacher, including train on teacher-generated text Kim and Rush (2016); Taori et al. (2023) and teacher output at each step Sanh et al. (2019). There are also many works Agarwal et al. (2023); Passalis and Tefas (2018); Wen et al. (2023) exploring the metric of distribution distance in KD. Work Huang et al. (2022) points out the inadequacy of KL distance when the difference between teacher and student distribution is large, and proposed a relaxed matching term as a supplement to KL divergence. Concurrent works Cho and Hariharan (2019); Mirzadeh et al. (2020) also proposed a similar phenomenon when the student and teacher model sizes significantly differ.
KD for LLM
has received attention in recent years due to the booming open source LLMs community Zhang et al. (2022a); Touvron et al. (2023). KD in large models is usually divided into black-box distillation Taori et al. (2023); Chiang et al. (2023) and white-box distillation Kim et al. (2024). The black-box KD takes the text generated by the teacher as knowledge and performs distillation, while the wite-box KD use the model intermediate layers or output logits distribution Jiao et al. (2020); Wang et al. (2020). Recent works have recognized the inadequacy of KL divergence in knowledge distillation of LLMs Gu et al. (2023) and applied reverse KL divergence to KD of LLMs. Concurrent work Wu et al. (2024) illustrates the shortcomings of reverse KLD and KLD and designs a KL divergence that adaptively allocates weights.
6 Conclusion
In this work we propose a novel method for knowledge distillation of LLMs from the perspective of direct preference learning. We introduce implicit reward and output preference models in knowledge distillation of LLMs and re-formulate the goal of KD. We perform theoretical derivation to obtain a new distillation framework. Experiments show that this framework performs well and is applicable to a wide range of models and data. In addition, we conducts extensive studies on the reward function, preference formula form, showing its importance and providing insights for subsequent work.
Limitations
We recognize several limitations of this work. Firstly, the form of preference will affect the situation of model learning $y_{t}$ and $y_{s}$. In some cases, direct preference learning will cause the model to focus too much on the relative probability of the two, rather than on the absolute probability of $y_{t}$ that we expect to obtain. As mentioned in Section 4.4, there are many works that explore and improve different forms of preference. This paper conducts preliminary experiments on these forms of preference and shows that some of the methods may be effective. However, in the field of instruction fine-tuning involved in this paper, the above direct preference situation is not serious. However, the effectiveness of other methods in the preliminary experiments of this paper shows that there may be other problems caused by preferences in the field of LLM KD, and the performance can be further improved through design methods. Section 4.4 has introduced preliminary experiments, and we hope that the DPKD methods and this part of the experiment can provide inspiration for future work.
Appendix A Complete Derivation of DPKD Objective
The optimization goal of first step is defined as:
$\max_{\theta}\mathbb{E}\,[\,r_{p}(y|x)-\beta\,\text{KLD}\,\,(q_{\theta}(y|x))\|p(y|x))\,]$ (17)
Transform Equation 17 as follows:
$\max_{\theta}\mathbb{E}\left[r(x,y)-\beta\log\frac{q_{\theta}(y\mid x)}{p(y\mid x)}\right]$ (18)
$=$ $\min_{\theta}\mathbb{E}\left[\log\frac{q_{\theta}(y\mid x)}{p(y\mid x)}-\frac{1}{\beta}r(x,y)\right]$
$=$ $\min_{\theta}\mathbb{E}\left[\log\frac{q_{\theta}(y\mid x)}{\frac{1}{Z(x)}p(y\mid x)\exp\left(\frac{1}{\beta}r(x,y)\right)}-\log Z(x)\right]$
$=$ $\min_{\theta}\mathbb{E}\left[\log\frac{q_{\theta}(y\mid x)}{q^{*}(y\mid x)}-\log Z(x)\right]$
where $q^{*}(y|x)\mathrel{\coloneqq}\frac{1}{Z(x)}p(y|x)\exp\left(\frac{1}{\beta}r(x,y)\right)$, and $Z\left(x\right)$ is the normalized function of the distribution, which is required to be independent of $\theta$ and $y$. It is defined as $Z(x)\mathrel{\coloneqq}\sum_{y}p(y|x)\exp\left(\frac{1}{\beta}r(x,y)\right)$.
Substituting into the optimization target, we can get the optimization target:
$\min_{\theta}\mathbb{E}_{x}\left[\mathbb{E}_{y}\left[\log\frac{q_{\theta}(y\mid x)}{q^{*}(y\mid x)}\right]-\log Z(x)\right]$ (19)
$=$ $\min_{\theta}\mathbb{E}_{x}\left[\text{KLD}\left(q_{\theta}(y\mid x)\|q^{*}(y\mid x)\right)-\log Z(x)\right]$
Since $Z\left(x\right)$ is a normalized function independent of $\theta$, the minimum value from the optimization objective  19 is achieved by minimizing the first KL divergence term. We can derive the optimal solution of Equation 19:
$q_{\theta}(y\mid x)=q^{*}(y\mid x)$ (20)
From Equation 20 we can obtain that:
$r^{*}(x,y)=\beta\log\frac{q^{*}(y\mid x)}{p(y\mid x)}+\beta\log Z(x)$ (21)
Given the same prompt x, the outputs of the student model and the teacher’s model are respectively denoted as $y_{s}$ and $y_{t}$. The purpose of KD is to fit the distribution of the student model to teacher model, which can also be understood as we expect the student model to have a greater probability of outputting results similar to the teacher model. From the BT model, we can obtain the following probability:
$p$ ${}^{*}\left(y_{t}\succ y_{s}\mid x\right)$ (22)
$=\sigma\left(\beta\log\frac{q_{\theta}\left(y_{t}\mid x\right)}{p\left(y_{t}\mid x\right)}-\beta\log\frac{q_{\theta}\left(y_{s}\mid x\right)}{p\left(y_{s}\mid x\right)}\right)$
Since our goal is to maximize the probability of model outputing $y_{t}$ rather than $y_{s}$, which is:
$\max_{\theta}\,\,\mathbb{E}\log\,p^{*}\left(y_{t}\succ y_{s}\mid x\right)$ (23)
$\min_{\theta}-\mathbb{E}\left[\log\sigma\left(\beta\log\frac{q_{\theta}\left(y_{t}\right)}{p\left(y_{t}\right)}-\beta\log\frac{q_{\theta}\left(y_{s}\right)}{p\left(y_{s}\right)}\right)\right]$ (24)
where $q_{\theta}\left(y_{t}\right)$ denotes $q_{\theta}\left(y_{t}\mid x\right)$ for short, and same for $p$ and $y_{s}$.
Appendix B Complete Derivation of DPKD Gradient
Based on our target definition, we can derive the gradient of DPKD through the basic chain rule and perform analysis. From Equation 8, we can perform variable substitution and define $u\mathrel{\coloneqq}\beta\log\frac{q_{\theta}\left(y_{s}\mid x\right)}{p\left(y_{s}\mid x\right)}-\beta\log\frac{q_{\theta}\left(y_{t}\mid x\right)}{p\left(y_{t}\mid x\right)}$, and the gradient of the loss can be expressed as
$\nabla_{\theta}\mathcal{L}$ $=-\nabla_{\theta}\mathbb{E}_{\left(x,y_{t},y_{s}\right)\sim\mathcal{D}}\left[\log\sigma\left(u\right)\right]$ (25)
$=-\mathbb{E}_{\left(x,y_{t},y_{s}\right)\sim\mathcal{D}}\left[\frac{\sigma^{\prime}(u)}{\sigma(u)}\nabla_{\theta}(u)\right]$
Define the variable u as follows:
$u=\beta\log\frac{q_{\theta}\left(y_{s}\mid x\right)}{p\left(y_{s}\mid x\right)}-\beta\log\frac{q_{\theta}\left(y_{t}\mid x\right)}{p\left(y_{t}\mid x\right)}$ (26)
Substitute and perform gradient deformation:
$\nabla_{\theta}\mathcal{L}_{\mathrm{DPKD}}\left(\theta\right)=-\mathbb{E}_{\left(x,y_{t},y_{s}\right)\sim\mathcal{D}}\left[\frac{\sigma^{\prime}(u)}{\sigma(u)}\nabla_{\theta}(u)\right]$ (27)
Since $\sigma^{\prime}(x)=\sigma(x)(1-\sigma(x))$, we can derive the complete form of gradient:
$\nabla_{\theta}\mathcal{L}=$ $-\mathbb{E}\left[\beta\sigma\left(\beta\log\frac{q_{\theta}\left(y_{t}\right)}{p\left(y_{t}\right)}-\beta\log\frac{q_{\theta}\left(y_{s}\right)}{p\left(y_{s}\right)}\right)\right.$ (28)
$\left.\left[\nabla_{\theta}\log q_{\theta}\left(y_{t}\right)-\nabla_{\theta}\log q_{\theta}\left(y_{s}\right)\right]\right]$
For further analysis, we make a representable approximation of the reward function form in Equation 21
$\hat{r}_{\theta}(x,y)=\beta\log\frac{q_{\theta}(y\mid x)}{p(y\mid x)}$ (29)
and substitute it into the Equation 28:
$\nabla_{\theta}\mathcal{L}=$ $-\beta\,\mathbb{E}\left[\sigma\left(\hat{r}_{\theta}(x,y_{t})-\hat{r}_{\theta}(x,y_{s})\right)\right.$ (30)
$\left.\left[\nabla_{\theta}\log q_{\theta}\left(y_{t}\right)-\nabla_{\theta}\log q_{\theta}\left(y_{s}\right)\right]\right]$
Appendix C Complete Result of Different Generation Lengths
Method $\left(0,30\right)$ $\left(30,70\right)$ $\left(70,\infty\right)$
KD 28.04 21.65 15.07
SeqKD 29.32 24.67 16.16
SFT 30.10 25.28 16.09
DPKD 30.98 26.69 17.26
Table 5: Exact value of RougeL with different ranges of generation lengths.
The complete result of Figure 4 is shown in Table 5. From the results, we can see that as the length of the golden response increases, the absolute value of RougeL decreases. This is because the difficulty of the corresponding task of the instruction increases. However, we can still see the advantage of DPKD, as well as the huge lead of DPKD in the middle range of length.